OpenStudy (anonymous):

Find the distance for the following points A ( -2, -3) B (-7, -8). If necessary round your answer to the nearest tenth

OpenStudy (anonymous):

use formula for the distance: d^2=(-2-(-7))^2+(-3-(-8))^2

OpenStudy (anonymous):

d^2=25+25=50 \[d=\sqrt{50}\]=... use calculator to compute